OpenStudy (anonymous):

what is the next number in the sequence: 3, 8, 18, 38...?

OpenStudy (anonymous):

It would be 2 times the number before plus 2 so 38(2)+2 78 78 would be it

OpenStudy (ddcamp):

Look at the differences between each term. Is there a clear pattern there?

OpenStudy (anonymous):

Yup, 3(2)+2=8 8(2)+2=18 18(2)+2=38 So 38(2)+2=78 @DDCamp